OpenStudy (anonymous):

find the equation of the parabola that asses through each set of ppoints (2,0)(3,-2)(1,-2)

OpenStudy (anonymous):

*****Passes opps!

OpenStudy (anonymous):

Is It Multiple Choice.?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay Cause iDon't Know Whether To Use y=(x-h)^2+k or y=ax^2+bx+c

OpenStudy (anonymous):

its y=ax^2+bx+c that im sure of!

hero (hero):

Use the one where you will be left with a system of two equations

hero (hero):

Insert the points into the y = (x - h)^2 + k

OpenStudy (anonymous):

Alright plug-in the first point o=a(2)^2+b(2)+c simplify 0=4a+2b+c plug-in the second point -2=a(3)^2+b(3)+c simplfy -2=9a+3b+c last point -2=a(1)^2+b(1)+c simplify -2=a+b+c

OpenStudy (anonymous):

Do You Have A Matrix Calcultor.?

OpenStudy (anonymous):

No i dont

OpenStudy (anonymous):

How would i insert in that equation @Hero

hero (hero):

0 = (2 - h)^2 + k -2 = (3 - h)^2 + k

OpenStudy (anonymous):

and -2=(1-h)^2+k

OpenStudy (anonymous):

then what? @Hero

OpenStudy (anonymous):

wait what happened

OpenStudy (anonymous):

@Pinky<3

OpenStudy (anonymous):

a=-2, b=8, and c=-8

OpenStudy (anonymous):

im Sorry iDeleted It

OpenStudy (anonymous):

where did you get that

OpenStudy (anonymous):

If You Use A Matrix Calculator It Helps http://www.bluebit.gr/matrix-calculator/linear_equations.aspx

hero (hero):

When I did it, h didn't come out to an even number so I got discouraged.

hero (hero):

But you should learn how to do it without a calculator

OpenStudy (anonymous):

yeah i want to do this without a calculator can you help

hero (hero):

Actually, now I know what went wrong; The formula is y = a(x - h)^2 + k

OpenStudy (anonymous):

ok

hero (hero):

You wrote y = (x - h)^2 + k the first time

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

how can i get this without a calculator

hero (hero):

The easiest way to do this is to realize that (3,-2) and (1,-2) are equidistant from the x coordinate of symmetry

hero (hero):

so x = 2 is the x coordinate of the vertex which is very helpful.

hero (hero):

Because they give us a point (2,0) which happens to be the vertex.

hero (hero):

So now all we have to do is plug in a given point and the vertex, then solve for a

OpenStudy (anonymous):

ohhh

hero (hero):

-2 = a(3 - 2)^2 + 0 -2 = a(1)^2 + 0 -2 = a + 0 -2 = a

hero (hero):

So the equation of the parabola should be y = -2(x - 2)^2

OpenStudy (anonymous):

my book says the answer is y=-2x^2+8x-8

hero (hero):

Expand -2(x - 2)^2 and see what you get

OpenStudy (anonymous):

how do i do that??

OpenStudy (anonymous):

how do i expand it!

hero (hero):

-2(x - 2)^2 = -2(x - 2)(x - 2) = -2(x(x- 2) - 2(x-2)) = -2(x^2 - 2x - 2x - 4) = -2(x^2 - 4x - 4) = -2x^2 + 8x + 8

OpenStudy (anonymous):

thanks you guys soooooo much!!!!

OpenStudy (anonymous):

thanks @Hero